# Biorthogonal ensembles of derivative type

### v1

@@ 1. Introduction
which is (see, e.g., [24, Section 3.2]) the eigenvalue distribution of the $N\times N$ Laguerre Unitary Ensemble (LUE), or Wishart-Laguerre Ensemble, of Hermitian positive-definite random matrices. Notice that for $\nu=0$, $w$ is precisely N1N-1 times differentiable in $L^{1}$-sense, with the last derivative being integrable but not continuous.
@@
While these are the two most classical examples of polynomial ensembles of additive derivative type, others occur as eigenvalue distributions of sums of independent random matrices [36, 14, 41]. The class of functions $w$ for which (1.6) defines a positive measure for every $N\in\mathbb{N}$ is rather restricted: it contains only those functions $w$ for which
@@
For $\tilde{w}(y)={y^{\mu}(1-y)^{\nu}\mathbbm{1}_{(0,1)}(y)}$ with $\mu>-1,\nu>N-1$, (1.10) is the eigenvalue distribution in the Jacobi Unitary Ensemble (JUE) [24, Section 3.6].
@@
If we set $\tilde{w}(y)=y^{\beta}(1+y)^{-\beta-\gamma-1}$ with $\beta>-1,\gamma>N-1$, (1.10) is the eigenvalue distribution in the Cauchy-Lorentz Unitary Ensemble (C-LUE), see [24, Section 2.5] and [54].
@@ 2.1. Double contour integral expression for the correlation kernel
where the contour $\mathcal{C}$ of $u$ is encircling 0 in the positive direction. The latter expression was obtained in [36, Corollary III.3] for the correlation kernel of any orthogonal polynomial ensemble of additive derivative type.
@@
Compared to the definition of Pólya frequency functions in [35], here we also require $w$ being integrable, and hence the monomial factor is omitted. We will prove in Proposition 9 that this function belongs to $\mathcal{V}_{c_{-},c_{+}}^{N}$ with $c_{-}=\max_{b_{j}<0}b_{j}^{-1}$ and $c_{+}=\min_{b_{j}>0}b_{j}^{-1}$ provided that $\tau>0$ or $\tau=0$ with at least $N$ non-zero $b_{j}$-values. If there are no positive $b_{j}$-values, we can take $c_{+}>0$ arbitrarily, and if there are no negative $b_{j}$-values, we can take $c_{-}<0$ arbitrarily. As the corresponding functions $w$ give rise to non-negative biorthogonal ensembles of derivative type, this provides us with a practical method to construct such ensembles. The special case $b_{j}=0$ for all $j$ is (up to a shift by $\gamma$ and scaling with $\tau$) the GUE with external source, while $\tau=0$, $b_{1}=\ldots=b_{N+\nu}=1$ and $b_{j}=0$ for $j>N+\nu$ is the LUE with external source (again, up to shifting and scaling), with an integer parameter $\nu$.
@@ 2.2. Limit kernels for additive perturbations of LUE
For our next result, we return to polynomial ensembles of additive derivative type as given in (1.6). They have the remarkable property of being closed under additive convolution. More precisely, suppose that $A_{1}$ and $A_{2}$ are independent $N\times N$ Hermitian random matrices whose eigenvalue distributions take the form (1.6) with $w=w_{1}$ for $A_{1}$ and $w=w_{2}$ for $A_{2}$, with $w_{1},w_{2}\in\mathcal{U}_{c_{-},c_{+}}^{N}$. Then, a special case of [36, Theorem II.4] states that the eigenvalue distribution of $A_{1}+A_{2}$ is again a polynomial ensemble of derivative type (1.6), now with $w$ being the (additive) convolution of $w_{1}$ and $w_{2}$,
@@
As we will explain in detail in Section 3.6, this fact in combination with Theorem 1 induces a simple transformation on the level of correlation kernels: the correlation kernels of $A_{1}$ and $A_{2}$ have the form (2.1) with $W=W_{1}$ and $W=W_{2}$ given in terms of $w=w_{1}$ and $w=w_{2}$ by (2.3), and the correlation kernel of $A_{1}+A_{2}$ has the form (2.1) with $W=W_{1}W_{2}$. A similar property holds for polynomial ensembles of multiplicative derivative type [37, Corollary 3.4].
@@
where $\Sigma$ is a simple positively oriented closed contour enclosing 0, and $c>0$ is such that $c+i\mathbb{R}$ lies at the right of $\Sigma$. This same expression for the correlation kernel was derived in [10, Equation (7.14)]. Another double contour integral expression was given in [14, Theorem 2.3] for the additive convolution of a polynomial ensemble with a GUE matrix, but it is less explicit in the sense that the integrand contains a correlation kernel of the LUE.
@@
$\mathbb{K}^{\rm pBessel}(x,x^{\prime};r):=\int_{\tilde{\Sigma}}\frac{\mathrm{d}s}{2\pi i}\int_{c+i\mathbb{R}}\frac{\mathrm{d}t}{2\pi i}\left(\frac{s}{t}\right)^{\nu}\frac{W(rt)}{W(rs)}\frac{\exp(\frac{1}{4t}-xt-\frac{1}{4s}+x^{\prime}s)}{t-s}.$ (2.16)
Here $c_{-}/r<c<0$, and $\tilde{\Sigma}$ is a positively oriented circle with center $c/3$ and radius $|c|/3$, passing through 0, see Figure 2.
@@
$\mathbb{K}^{\rm pBessel}(x,x^{\prime};0):=\int_{\tilde{\Sigma}}\frac{\mathrm{d}s}{2\pi i}\int_{c+i\mathbb{R}}\frac{\mathrm{d}t}{2\pi i}\left(\frac{s}{t}\right)^{\nu}\frac{\exp(\frac{1}{4t}-xt-\frac{1}{4s}+x^{\prime}s)}{t-s}.$
This is actually a known double contour integral expression for the hard edge Bessel kernel, see [31].
@@
$\lim_{N\to\infty}\frac{1}{N}K_{N}^{\rm LUE+GUE}\left(\frac{x}{4N},\frac{x^{\prime}}{4N};\frac{r}{4N}\right)=\int_{\tilde{\Sigma}}\frac{\mathrm{d}s}{2\pi i}\int_{c+i\mathbb{R}}\frac{\mathrm{d}t}{2\pi i}\left(\frac{s}{t}\right)^{\nu}\frac{\exp(\frac{r^{2}t^{2}}{2}+\frac{1}{4t}-xt+\frac{r^{2}s^{2}}{2}-\frac{1}{4s}+x^{\prime}s)}{t-s}.$ (2.17)
This is an alternative (and simpler) expression of a limit kernel that appeared in [12, Eqn. (2.9)].
@@ 2.3. Limit kernels for Muttalib-Borodin deformations
  • For $\theta=1$, $\eta=0$ and $W(z)={\prod_{j=1}^{M}\Gamma(z+\nu_{j})}$, (2.19) is the singular value distribution of products of Ginibre matrices, and the kernel (2.25) is equivalent to the Meijer-G kernel; see [43, Eqn. (5.7)] and [13, Eqn. (1.11)–(1.12) and Appendix A].
  • For $\eta=0$, $\theta>1$ and $W(z)={\Gamma(z+\nu)}$, (2.19) reduces to the Laguerre Muttalib-Borodin ensembles, and the kernel (2.25) is equal to Wright’s generalized Bessel kernel, see [13, Eqn. (1.11)–(1.14) and Appendix A]. Note that the condition $\theta>1$ is needed for (2.22) to hold.
@@ 3.3. Correlation kernel
It was proved in [10, Proposition 2.1] that (3.13) satisfies the biorthogonality relations (1.5), and that the correlation kernel defined by (1.4) has the double contour integral expression (2.1).
@@ 3.4. The confluent case
This is a confluent Vandermonde determinant; see a recent review [44] and also [24, Eqn. (1.174)]. It was shown in [1] that
@@ 3.6. Additive convolutions
We now give more details about additive convolutions of polynomial ensembles of derivative type, as announced in Section 2.2. Let $A,B$ be independent Hermitian $N\times N$ random matrices with eigenvalue distribution of $A$ the polynomial ensemble of derivative type (1.11) associated to $w_{1}\in\mathcal{U}_{c_{-},c_{+}}^{N}$ and eigenvalue distribution of $B$ the polynomial ensemble of derivative type (1.11) associated to $w_{2}\in\mathcal{U}_{c_{-},c_{+}}^{N}$. By [36, Thm II.4] the eigenvalue distribution of the random matrix $X=A+B$ is the polynomial ensemble (1.6) associated to functions
@@ 4.2. LUE with additive Pólya perturbation
$\frac{1}{4N}K_{N}^{\rm pLUE}\left(\frac{x}{4N},\frac{x^{\prime}}{4N};\frac{r}{4N}\right)=e^{-\frac{x-x^{\prime}}{8N}}\oint_{\tilde{\Sigma}}\frac{\mathrm{d}s}{2\pi i}\int_{\tilde{c}+i\mathbb{R}}\frac{\mathrm{d}t}{2\pi i}\left(\frac{\frac{1}{8N}-s}{\frac{1}{8N}-t}\right)^{\nu}\frac{W\left(rt+\frac{r}{8N}\right)}{W\left(rs+\frac{r}{8N}\right)}\\ \times\quad\frac{\exp\left(N\log\frac{1+\frac{1}{8Nt}}{1-\frac{1}{8Nt}}-xt\right)}{\exp\left(N\log\frac{1+\frac{1}{8Ns}}{1-\frac{1}{8Ns}}-x^{\prime}s\right)}\frac{1}{t-s}.$ (4.4)
@@
$\lim_{N\to\infty}\frac{1}{4N}K_{N}^{\rm pLUE}\left(\frac{x}{4N},\frac{x^{\prime}}{4N};\frac{r}{4N}\right)=\int_{\tilde{\Sigma}}\frac{\mathrm{d}s}{2\pi i}\int_{\tilde{c}+i\mathbb{R}}\frac{\mathrm{d}t}{2\pi i}\left(\frac{s}{t}\right)^{\nu}\frac{W\left(rt\right)}{W\left(rs\right)}\frac{\exp\left(\frac{1}{4t}-xt\right)}{\exp\left(\frac{1}{4s}-x^{\prime}s\right)}\frac{1}{t-s}.$

### v2

@@ 1. Introduction
which is (see, e.g., [27, Section 3.2]) the eigenvalue distribution of the $N\times N$ Laguerre Unitary Ensemble (LUE), or Wishart-Laguerre Ensemble, of Hermitian positive-definite random matrices. Notice that for $\nu=0$, $w$ is precisely N1N-1 times differentiable in $L^{1}$-sense, with the last derivative being integrable but not continuous.
@@
While these are the two most classical examples of polynomial ensembles of additive derivative type, others occur as eigenvalue distributions of sums of independent random matrices [39, 17, 44] and in dynamical random matrix ensembles [5, 6]. The class of functions $w$ for which (1.6) defines a positive measure for every $N\in\mathbb{N}$ is rather restricted: it contains only those functions $w$ for which
@@
For $\tilde{w}(y)={y^{\mu}(1-y)^{\nu}\mathbbm{1}_{(0,1)}(y)}$ with $\mu>-1,\nu>N-1$, (1.10) is the eigenvalue distribution in the Jacobi Unitary Ensemble (JUE) [27, Section 3.6].
@@
If we set $\tilde{w}(y)=y^{\beta}(1+y)^{-\beta-\gamma-1}$ with $\beta>-1,\gamma>N-1$, (1.10) is the eigenvalue distribution in the Cauchy-Lorentz Unitary Ensemble (C-LUE), see [27, Section 2.5] and [60].
@@ 2.1. Double contour integral expression for the correlation kernel
where the contour $\mathcal{C}$ of $u$ is encircling 0 in the positive direction. The latter expression was obtained in [39, Corollary III.3] for the correlation kernel of any orthogonal polynomial ensemble of additive derivative type.
@@
Remark 6.
The double contour integral expression (2.1) is similar to a double contour integral expression for the correlation kernel of Schur measures [55, 56]. In the latter, both integrations are over closed curves in the complex plane, which is related to the fact that the Schur measures live on a discrete lattice rather than on the real line. In [10, Section 4], certain continuous versions of Schur measures were studied, with correlation kernels of the same form as (2.1), but in the special case where $1/W$ is a polynomial.
@@
Note that the definition of Pólya frequency functions in [38] looks slightly different from what we have, namely they have the more general form
$W(z)=\mathcal{F}[w](-iz)=cz^{r}e^{\tau z^{2}/2+\gamma z}\prod_{j=1}^{\infty}\frac{e^{-b_{j}z}}{1-b_{j}z}.$ (2.12)
However, we require $w$ to be integrable, such that $W(0)=\int_{-\infty}^{+\infty}w(x)\mathrm{d}x\in(0,+\infty)$, which is only possible if $r=0$. We will prove in Proposition 9 that functions of the form (2.11) belong to $\mathcal{V}_{c_{-},c_{+}}^{N}$ with $c_{-}=\max_{b_{j}<0}b_{j}^{-1}$ and $c_{+}=\min_{b_{j}>0}b_{j}^{-1}$ provided that $\tau>0$ or $\tau=0$ with at least $N$ non-zero $b_{j}$-values. If there are no positive $b_{j}$-values, we can take $c_{+}>0$ arbitrarily, and if there are no negative $b_{j}$-values, we can take $c_{-}<0$ arbitrarily. As the corresponding functions $w$ give rise to non-negative biorthogonal ensembles of derivative type, this provides us with a practical method to construct such ensembles. The special case $b_{j}=0$ for all $j$ is (up to a shift by $\gamma$ and scaling with $\tau$) the GUE with external source, while $\tau=0$, $b_{1}=\ldots=b_{N+\nu}=1$ and $b_{j}=0$ for $j>N+\nu$ is the LUE with external source (again, up to shifting and scaling), with an integer parameter $\nu$.
@@ 2.2. Limit kernels for additive perturbations of LUE
For our next result, we return to polynomial ensembles of additive derivative type as given in (1.6). They have the remarkable property of being closed under additive convolution. More precisely, suppose that $A_{1}$ and $A_{2}$ are independent $N\times N$ Hermitian random matrices whose eigenvalue distributions take the form (1.6) with $w=w_{1}$ for $A_{1}$ and $w=w_{2}$ for $A_{2}$, with $w_{1},w_{2}\in\mathcal{U}_{c_{-},c_{+}}^{N}$. Then, a special case of [39, Theorem II.4] states that the eigenvalue distribution of $A_{1}+A_{2}$ is again a polynomial ensemble of derivative type (1.6), now with $w$ being the (additive) convolution of $w_{1}$ and $w_{2}$,
@@
As we will explain in detail in Section 3.6, this fact in combination with Theorem 1 induces a simple transformation on the level of correlation kernels: the correlation kernels of $A_{1}$ and $A_{2}$ have the form (2.1) with $W=W_{1}$ and $W=W_{2}$ given in terms of $w=w_{1}$ and $w=w_{2}$ by (2.3), and the correlation kernel of $A_{1}+A_{2}$ has the form (2.1) with $W=W_{1}W_{2}$. A similar property holds for polynomial ensembles of multiplicative derivative type [40, Corollary 3.4].
@@
where $\Sigma$ is a simple positively oriented closed contour enclosing 0, and $c>0$ is such that $c+i\mathbb{R}$ lies at the right of $\Sigma$. This same expression for the correlation kernel was derived in [13, Equation (7.14)]. Another double contour integral expression was given in [17, Theorem 2.3] for the additive convolution of a polynomial ensemble with a GUE matrix, but it is less explicit in the sense that the integrand contains a correlation kernel of the LUE.
@@
$\mathbb{K}^{\rm pBessel}(x,x^{\prime};r):=\int_{\tilde{\Sigma}}\frac{\mathrm{d}s}{2\pi i}\int_{c+i\mathbb{R}}\frac{\mathrm{d}t}{2\pi i}\frac{(-s)^{\nu}}{(-t)^{\nu}}\frac{W(rt)}{W(rs)}\frac{\exp(\frac{1}{4t}-xt-\frac{1}{4s}+x^{\prime}s)}{t-s}.$ (2.17)
Here $c_{-}/r<c<0$, and $\tilde{\Sigma}$ is a positively oriented circle with center $c/3$ and radius $|c|/3$, passing through 0, see Figure 2. The branch cut of the $\nu$-th power function is taken to be the negative real line so that both $s$ and $t$ have a branch cut at the positive real line, which does not enclose either of the contours.
@@
$\mathbb{K}^{\rm pBessel}(x,x^{\prime};0):=\int_{\tilde{\Sigma}}\frac{\mathrm{d}s}{2\pi i}\int_{c+i\mathbb{R}}\frac{\mathrm{d}t}{2\pi i}\frac{(-s)^{\nu}}{(-t)^{\nu}}\frac{\exp(\frac{1}{4t}-xt-\frac{1}{4s}+x^{\prime}s)}{t-s}.$ (2.18)
This is actually a known double contour integral expression for the hard edge Bessel kernel; see [34] and Appendix A.
@@
$\lim_{N\to\infty}\frac{1}{N}K_{N}^{\rm LUE+GUE}\left(\frac{x}{4N},\frac{x^{\prime}}{4N};\frac{r}{4N}\right)=\int_{\tilde{\Sigma}}\frac{\mathrm{d}s}{2\pi i}\int_{c+i\mathbb{R}}\frac{\mathrm{d}t}{2\pi i}\frac{(-s)^{\nu}}{(-t)^{\nu}}\frac{\exp(\frac{r^{2}t^{2}}{2}+\frac{1}{4t}-xt+\frac{r^{2}s^{2}}{2}-\frac{1}{4s}+x^{\prime}s)}{t-s}.$ (2.19)
This is an alternative (and simpler) expression of a limit kernel computed in a different way in [15, Eqn. (2.9)]. The same limit kernel appeared in [57, Formula (1.9)] in the context of a first passage percolation problem.
@@ 2.3. Limit kernels for Muttalib-Borodin deformations
  • For $\theta=1$, $\eta=0$ and $W(z)={\prod_{j=1}^{M}\Gamma(z+\nu_{j})}$, (2.21) is the singular value distribution of products of Ginibre matrices, and the kernel (2.27) is equivalent to the Meijer-G kernel; see [46, Eqn. (5.7)] and [16, Eqn. (1.11)–(1.12) and Appendix A].
  • For $\eta=0$, $\theta>1$ and $W(z)={\Gamma(z+\nu)}$, (2.21) reduces to the Laguerre Muttalib-Borodin ensembles, and the kernel (2.27) is equal to Wright’s generalized Bessel kernel, see [16, Eqn. (1.11)–(1.14) and Appendix A]. Note that the condition $\theta>1$ is needed for (2.24) to hold.
@@ 3.3. Correlation kernel
It was proved in [13, Proposition 2.1] that (3.13) satisfies the biorthogonality relations (1.5), and that the correlation kernel defined by (1.4) has the double contour integral expression (2.1).
@@ 3.4. The confluent case
This is a confluent Vandermonde determinant; see a recent review [47] and also [27, Eqn. (1.174)]. It was shown in [1] that
@@ 3.6. Additive convolutions
We now give more details about additive convolutions of polynomial ensembles of derivative type, as announced in Section 2.2. Let $A,B$ be independent Hermitian $N\times N$ random matrices with eigenvalue distribution of $A$ the polynomial ensemble of derivative type (1.11) associated to $w_{1}\in\mathcal{U}_{c_{-},c_{+}}^{N}$ and eigenvalue distribution of $B$ the polynomial ensemble of derivative type (1.11) associated to $w_{2}\in\mathcal{U}_{c_{-},c_{+}}^{N}$. By [39, Thm II.4] the eigenvalue distribution of the random matrix $X=A+B$ is the polynomial ensemble (1.6) associated to functions
@@ 4.2. LUE with additive Pólya perturbation
$\frac{1}{4N}K_{N}^{\rm pLUE}\left(\frac{x}{4N},\frac{x^{\prime}}{4N};\frac{r}{4N}\right)=e^{-\frac{x-x^{\prime}}{8N}}\oint_{\tilde{\Sigma}}\frac{\mathrm{d}s}{2\pi i}\int_{\tilde{c}+i\mathbb{R}}\frac{\mathrm{d}t}{2\pi i}\frac{(\frac{1}{8N}-s)^{\nu}}{(\frac{1}{8N}-t)^{\nu}}\frac{W\left(rt+\frac{r}{8N}\right)}{W\left(rs+\frac{r}{8N}\right)}\\ \times\quad\frac{\exp\left(N\log\frac{1+\frac{1}{8Nt}}{1-\frac{1}{8Nt}}-xt\right)}{\exp\left(N\log\frac{1+\frac{1}{8Ns}}{1-\frac{1}{8Ns}}-x^{\prime}s\right)}\frac{1}{t-s}.$ (4.4)
@@
$\lim_{N\to\infty}\frac{1}{4N}K_{N}^{\rm pLUE}\left(\frac{x}{4N},\frac{x^{\prime}}{4N};\frac{r}{4N}\right)=\int_{\tilde{\Sigma}}\frac{\mathrm{d}s}{2\pi i}\int_{\tilde{c}+i\mathbb{R}}\frac{\mathrm{d}t}{2\pi i}\frac{(-s)^{\nu}}{(-t)^{\nu}}\frac{W\left(rt\right)}{W\left(rs\right)}\frac{\exp\left(\frac{1}{4t}-xt\right)}{\exp\left(\frac{1}{4s}-x^{\prime}s\right)}\frac{1}{t-s}.$
Here we choose arguments between $\pi$ and $\pi$, such that both $(-s)^{\nu}$ and $(-t)^{\nu}$ have the same branch cut on $0,+\infty)$. The contour $\tilde{c}+i\mathbb{R}$ does not intersect with the branch cut, while $\tilde{\Sigma}$ approaches the branch point 0 from the vertical direction.
@@ 4.3. Muttalib-Borodin type biorthogonal ensembles of derivative type
Appendix A Different expressions for the hard edge Bessel kernel
The standard expressions for the hard edge Bessel kernel, see e.g. [27], are
$K_{\nu}(x,x^{\prime})=\frac{\sqrt{x^{\prime}}\,J_{\nu}(\sqrt{x})J_{\nu}^{\prime}(\sqrt{x^{\prime}})-\sqrt{x}\,J_{\nu}(\sqrt{x^{\prime}})J_{\nu}^{\prime}(\sqrt{x})}{2(x-x^{\prime})},\qquad\nu>-1,\quad x,x^{\prime}>0,$ (A.1)
and
$K_{\nu}(x,x^{\prime})=\frac{1}{4}\int_{0}^{1}J_{\nu}(\sqrt{xr})J_{\nu}(\sqrt{x^{\prime}r})\mathrm{d}r,\qquad\nu>-1,\quad x,x^{\prime}>0.$ (A.2)
Here $J_{\nu}$ denotes the Bessel function of the first kind, which can be expressed as a contour integral known as Schäfli’s integral [25, Formula (10.9.19)],
$J_{\nu}(\sqrt{z})=\frac{1}{2\pi i}\left(\frac{z}{4}\right)^{\nu/2}\int_{\gamma}v^{-\nu-1}e^{v-\frac{z}{4v}}\,\mathrm{d}v,\qquad\nu>-1,\ z>0,$ (A.3)
where $\gamma$ encloses the branch cut $(-\infty,0]$ in the counterclockwise direction. Note that for integer $\nu$, the integrand does not have a branch cut, such that one may equivalently take $\gamma$ to be a closed loop around 0. We will see that our expression for the Bessel kernel (2.18) is equivalent to (A.2) for $\nu\geq 0$.
Setting $u=\frac{z}{4v}$ in (A.3), we obtain
$J_{\nu}(\sqrt{z})=-\frac{1}{2\pi i}\left(\frac{z}{4}\right)^{-\nu/2}\oint_{\Sigma}u^{\nu-1}e^{-u+\frac{z}{4u}}\mathrm{d}u,$ (A.4)
where $\Sigma$ is the image of $\gamma$ under the map $v\mapsto\frac{z}{4v}$ (for $z>0$). In other words, it is a closed loop going through 0 and not intersecting $(-\infty,0)$, oriented clockwise and smooth except at 0. This integral is absolutely integrable for $\nu>-1$ if $\Sigma$ approaches the essential singularity 0 with angles $\pi\pm\theta$, $0\leq\theta<\pi$.
Substituting (A.3) and (A.4) into (A.2) and changing variables v=xtv=-xt, $u=-x^{\prime}s$ (with $x,x^{\prime}>0$), we obtain
$K_{\nu}(x,x^{\prime})=-\frac{1}{4}\left(\frac{x^{\prime}}{x}\right)^{\nu/2}\int_{0}^{1}\mathrm{d}r\left(\frac{1}{2\pi i}\int_{\tilde{\gamma}}(-t)^{-\nu-1}e^{-xt+\frac{r}{4t}}\mathrm{d}t\right)\left(\frac{1}{2\pi i}\oint_{\tilde{\Sigma}}(-s)^{\nu-1}e^{x^{\prime}s-\frac{r}{4s}}\mathrm{d}s\right),$ (A.5)
where $\tilde{\gamma}$ encloses $0,+\infty)$ with clockwise orientation, and $\tilde{\Sigma}$ is a closed positively oriented loop going through 0 and not intersecting $(0,+\infty)$, smooth except at 0 and approaching 0 with angles $\pm\theta$, $0\leq\theta<\pi$ (note that we have changed the orientation of both contours). We moreover choose $\tilde{\gamma}$ and $\tilde{\Sigma}$ such that they do not intersect. By Fubini’s theorem, we can bring the $r$-integral inside and compute it explicitly. We then obtain
$K_{\nu}(x,x^{\prime})=\left(\frac{x^{\prime}}{x}\right)^{\nu/2}\int_{\tilde{\gamma}}\frac{\mathrm{d}t}{2\pi i}\oint_{\tilde{\Sigma}}\frac{\mathrm{d}s}{2\pi i}\frac{(-s)^{\nu}}{(-t)^{\nu}}\frac{e^{-xt}}{e^{-x^{\prime}s}}\frac{e^{\frac{1}{4t}-\frac{1}{4s}}}{t-s}-\left(\frac{x^{\prime}}{x}\right)^{\nu/2}\int_{\tilde{\gamma}}\frac{\mathrm{d}t}{2\pi i}\oint_{\tilde{\Sigma}}\frac{\mathrm{d}s}{2\pi i}\frac{(-s)^{\nu}}{(-t)^{\nu}}\frac{e^{-xt}}{e^{-x^{\prime}s}}\frac{1}{t-s}.$
The second term in the above expression vanishes by Cauchy’s theorem because the $s$-integrand is analytic inside $\tilde{\Sigma}$ and continuous up to the boundary. The first term does not vanish because the $s$-integrand is unbounded inside $\tilde{\Sigma}$ near the essential singularity at 0. We finally obtain
$K_{\nu}(x,x^{\prime})=\left(\frac{x^{\prime}}{x}\right)^{\nu/2}\int_{\tilde{\gamma}}\frac{\mathrm{d}t}{2\pi i}\oint_{\tilde{\Sigma}}\frac{\mathrm{d}s}{2\pi i}\frac{(-s)^{\nu}}{(-t)^{\nu}}\frac{e^{-xt+\frac{1}{4t}}}{e^{-x^{\prime}s+\frac{1}{4s}}}\frac{1}{t-s}.$ (A.6)
We also observe that the $t$-integrand is $\mathcal{O}(|t|^{-\nu-1})$ as $t\to\infty$ in the half plane $\mathrm{Re}\,z\geq c$ for $c>0$, such that we can deform the contour $\tilde{\gamma}$ to the vertical line $c+i\mathbb{R}$. Similarly, the $s$-integrand is $\mathcal{O}(|s|^{\nu})$ as $s\to 0$ with $\mathrm{Re}\,s>0$, such that we can deform $\tilde{\Sigma}$ to a curve passing vertically through 0. Up to the cocycle factor $\left(\frac{x^{\prime}}{x}\right)^{\nu/2}$, this yields precisely our expression (2.18).